OpenStudy (anonymous):

Anna has leaned a ladder against the side of her house. If the ladder reaches up to a point that is 6 meters high and the bottom of the ladder is 2 meters from the wall, what angle does the ladder make with the ground? 12 degrees 18 degrees 52 degrees 72 degrees

OpenStudy (anonymous):

I'll be your best friend if you help me? LOL

OpenStudy (anonymous):

@hhelpplzzzz

OpenStudy (thomas5267):

Pythagoras theorem. 

OpenStudy (thomas5267):

I will assume the wall if not slanted.

OpenStudy (anonymous):

Why is the 2m longer than the 6?

OpenStudy (anonymous):

Its 72 I believe

OpenStudy (anonymous):

2^2+6^2=c^2 C=sqrt(40) Sin(x)={6/sqrt(40)}

OpenStudy (anonymous):

Take the inverse sine of your answer with a calculator.